157=6x^2/5 -233 what is x?
Answer by stanbon(75887)

157=6x^2/5 -233 what is x?
---
6x^(2/5) = 390
----
x^(2/5) = 65
---
[x^(2/5)]^(5/2) = 65^(6/5)
----
x = 149.79...
Cheers,
Stan H.